OpenStudy (anonymous):

G(x)=√(4-x) Evaluate the function : G(4x-x^2 ) help guys

OpenStudy (anonymous):

guys help

OpenStudy (anonymous):

cant sorry a little advanced.

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

just put '4x-x^2' in place of 'x' in G(x) what u get ?

OpenStudy (anonymous):

\[\sqrt{4-(4x-x^2)}\]

hartnn (hartnn):

correct, now simplify the quantity under sqrt sign.... and note that you get a perfect square inside!

OpenStudy (anonymous):

yeah like this \[\sqrt{4-4x+x^2}\] after that I don't know

hartnn (hartnn):

ok, so lets try to factor x^2-4x+4 tell me 2 numbers whose sum is -4 and product =4 ?

OpenStudy (anonymous):

you mean like this (x-2)(x-2)

hartnn (hartnn):

absolutely! thats \(\Large (x-2)^2\)

hartnn (hartnn):

so, \(\Large \sqrt{a^2}=|a|\) what about \(\Large \sqrt{(x-2)^2}=... ?\)

OpenStudy (anonymous):

x-2

hartnn (hartnn):

yes, more precisely its \(\Large |x-2|\). :)

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

and this is the final answer

hartnn (hartnn):

yup

hartnn (hartnn):

G(4x-x^2 ) = |x-2|

OpenStudy (anonymous):

sir I have another one I'm just stuck in math method would you mind helping me sir